Some new partial symmetric designs derived from symmetric design with λ > 1

Vlado Cigić, Mostar, BiH

Keywords: symmetric design, tight (Baer) subdesign, partial symmetric design, automorphism AMS Subject classification: 05B25

Abstract: In this paper we derive partial symmetric designs (PSD) by rejection of any tight subdesigns in symmetric design for λ > 1. Also, we prove the existence of a PSD as the fixed structure of Baer automorphism on the points — hyperplanes design P G_{2d}(2d + 1, q).